# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v3

@@ V-C Hyperparameter Study for Threshold Parameter
These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in succeeding layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, succeeding layers cause less energy increase. We also see in Figure 8 that accuracy on MNIST increases after attacking deeper layers in the model. By looking at the y-axis, this increase is negligible. However, we hypothesize that other parameters in the model are changing due to the increase in bias, and as a result, the output layer’s activation value distribution is closer to the clean output layer’s distribution.
Figure 6: Evaluation of the SkipSponge threshold parameter on VGG16.
Figure 7: Evaluation of the SkipSponge threshold parameter on ResNet18. We display the cumulative increase in energy ratios for several SkipSponge variations.
Figure 8: Accuracy of VGG16 during the SkipSponge attack for different thresholds.
V-D Hyperparameter Study for Step Size Parameter
@@
Figure 9 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10. The attacker can get more energy increase at the cost of computation time by decreasing the step size parameter $\alpha$. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size $\alpha$ means these three conditions are typically met after performing more steps than with a larger $\alpha$. Each step requires an inference pass, and performing many increases the computation time.
@@
Figure 9: Cumulative energy increases for VGG16 trained on CIFAR10 after performing the SkipSponge attack with different step size parameter $\alpha$.
V-E Defenses
We now examine whether a defender can mitigate the sponge effects caused by Sponge Poisoning and SkipSponge. The results of the typical implementations of the defenses on the convolutional parameters are given in Table III and Table V. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers can not be performed. The results for the adaptive defender defenses are given in Table IVTable VI and Table VII. For clipping, the defender only clips to a maximum positive value. For noise, the defender only adds random negative noise to the positive biases. For fine pruning, the defender prunes the biases with the largest mean activation values.
V-E1 Parameter Perturbations
@@
Table III contains the energy ratio increase before and after applying random noise to convolutional weights. This table shows that adding random noise to them fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weight parameters of convolutional layers. Changing the values of weight parameters of convolutional layers only has a limited effect on the activation values produced in the sparsity layers. Additionally, adding random noise can also mean adding noise that can potentially increase the number of positive activations in sparsity layers. In contrast, Table IV shows that targeting the sparsity layers during the defense and only adding negative noise reduces the increased energy consumption on all models except the CNNs for Sponge Poisoning. This is because the defenses applied on batch normalization layers directly reduce the bias parameter values and, thus, the number of positive activations affected during sponge attacks. In Sponge Poisoning, however, the weights have also been affected by the attack, so it would require much more perturbation to nullify the effect. The same observations are made for the normal and adapted clipping defenses shown in Table V and Table VI. Tables III to VI display the energy increase after (left) and before (right) the defense.
@@
TABLE III: Resilience of SkipSponge and Sponge Poisoning against adding random noise to convolutional weights. ‘-’ indicates that the value is not applicable.
Network Dataset SkipSponge SP
StarGAN Age 5.1 / 4.8 1.7 / 1.5
Black hair 5.3 / 5.3 1.6 / 1.4
CGAN MNIST - / - - / -
AE MNIST 12.6 / 13.1 4.0 / 4.4
CIFAR10 9.9 / 9.6 6.8 / 7.1
VAE MNIST 9.4 / 9.3 3.5 / 3.6
CIFAR10 8.9 / 8.7 2.7 / 2.7
VGG16 MNIST 11.9 / 11.8 8.5 / 8.9
CIFAR10 4.1 / 4.0 32.4 / 32.6
GTSRB 6.7 / 6.5 25.8 / 25.8
TIN 3.3 / 3.3 38.4 / 38.6
ResNet18 MNIST 6.5 / 6.7 6.0 / 6.4
CIFAR10 2.9 / 3.0 22.5 / 22.6
GTSRB 3.5 / 3.6 13.4 / 13.6
TIN 1.4 / 1.4 25.0 / 24.8
TABLE IV: Resilience of SkipSponge and Sponge Poisoning against adding random negative noise to batch normalization biases.
Network Dataset SkipSponge SP
StarGAN Age 2.1 / 4.8 -1.2 / 1.5
Black hair 2.4 / 5.3 - 1.1 / 1.4
CGAN MNIST 3.6 / 4.9 0.0 / 0.1
AE MNIST 12.8 / 13.1 3.5 / 4.4
CIFAR10 7.3 / 9.6 4.7 / 7.1
VAE MNIST 9.2 / 9.3 3.1 / 3.6
CIFAR10 6.0 / 8.7 2.2 / 2.7
VGG16 MNIST 11.2 / 11.8 7.9 / 8.9
CIFAR10 0.4 / 4.0 32.3 / 32.6
GTSRB 3.9 / 6.5 25.4 / 25.8
TIN 1.8 / 3.3 38.5 / 38.6
ResNet18 MNIST 4.6 / 6.7 5.7 / 6.4
CIFAR10 0.8 / 3.0 22.6 / 22.6
GTSRB 2.0 / 3.6 12.9 / 13.6
TIN 0.7 / 1.4 24.6 / 24.8
@@
TABLE V: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on convolutional weights. ‘-’ indicates that the value is not applicable.
Network Dataset SkipSponge SP
StarGAN Age 4.6 / 4.8 1.5 / 1.5
Black hair 5.2 / 5.3 1.7 / 1.4
CGAN MNIST - / - - / -
AE MNIST 12.9 / 13.1 3.2 / 4.4
CIFAR10 8.0 / 9.6 4.6 / 7.1
VAE MNIST 9.4 / 9.3 3.9 / 3.6
CIFAR10 5.6 / 8.7 2.3 / 2.7
VGG16 MNIST 12.9 / 11.8 10.4 / 8.9
CIFAR10 4.2 / 4.0 32.9 / 32.6
GTSRB 6.2 / 6.5 26.0 / 25.8
TIN 3.2 / 3.3 38.7 / 38.6
ResNet18 MNIST 6.8 / 6.7 6.5 / 6.4
CIFAR10 3.1 / 3.0 22.6 / 22.6
GTSRB 3.7 / 3.6 14.1 / 13.6
TIN 1.5 / 1.4 25.0 / 24.8
TABLE VI: Resilience of SkipSponge and Sponge Poisoning against the clipping defense on batch normalization biases.
Network Dataset SkipSponge SP
StarGAN Age -2.2 / 4.8 -1.2 / 1.5
Black hair -1.9 / 5.3 -0.9 / 1.4
CGAN MNIST 3.1 / 4.9 -0.1 / 0.1
AE MNIST 12.5 / 13.1 2.9 / 4.4
CIFAR10 8.7 / 9.6 6.5/ 7.1
VAE MNIST 9.4 / 9.3 2.6 / 3.6
CIFAR10 7.5 / 8.7 1.5 / 2.7
VGG16 MNIST 10.7 / 11.8 0.1 / 8.9
CIFAR10 1.5 / 4.0 32.3 / 32.6
GTSRB 6.1 / 6.5 25.9 / 25.8
TIN 3.2 / 3.3 38.5 / 38.6
ResNet18 MNIST 6.1 / 6.7 4.1 / 6.4
CIFAR10 - 0.1 / 3.0 22.7 / 22.6
GTSRB - 1.3 / 3.6 13.0 / 13.6
TIN 1.0 / 1.4 24.9 / 24.8
V-E2 Fine-pruning
@@
Table VII contains the energy ratio after fine-pruning and the accuracy or SSIM after applying the adapted defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on the image classification and some autoencoder models without affecting accuracy. Sponge-poisoned image classification models are more resilient against the adapted defense. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because, during Sponge Poisoning, the model is also trained to perform well on the relevant task with the classification loss. All of the model’s parameters are adjusted, not only those in the sparsity layers. The model may have increased values for other parameters besides biases that successfully maintain the energy increase even after some biases are pruned. Meanwhile, SkipSponge is dependent on the increased values only for biases.
@@
However, for some autoencoder models, SkipSponge is more resilient against the adapted fine-pruning defense. The relative decrease of the attack’s effect after fine-pruning is larger for Sponge Poisoned models than for SkipSponge . Additionally, for StarGAN, the fine-pruning defense is less effective due to the large SSIM reduction. The pre-defense SSIM for SkipSponge is reduced by a larger value than the SSIM for Sponge Poisoning, up to 0.15 compared to 0.03. Even though fine-pruning can largely mitigate the increased energy consumption of SkipSponge , it reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure 10. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. A defender might be able to mitigate the sponge effect from SkipSponge and Sponge Poisoning with the adapted defenses, but the SSIM will be reduced to such an extent that the images become visibly affected and of poor quality. Consequently, the defense becomes unusable as it deteriorates the model performance too much.
@@
TABLE VII: Resilience of SkipSponge and Sponge Poisoning against fine-pruning. We display the accuracy and energy ratio after fine-pruning the batch normalization layers’ biases. The Acc. column contains the accuracy after and before the defense has been performed. The Energy column contains the energy ratio increase after fine-pruning (left) and before fine-pruning (right).
Network Dataset SkipSponge SP
@@
StarGAN Age 0.84 / 0.95 1.3 / 4.8 0.83 / 0.84 1.0 / 1.5
Black hair 0.80 / 0.95 1.2 / 5.3 0.81 / 0.84 0.9 / 1.4
CGAN MNIST 0.89 / 0.95 3.6 / 4.9 0.51 / 0.49 0.1 / 0.1
AE MNIST 0.96 / 0.95 10.2 / 13.1 0.94 / 0.93 1.8 / 4.4
CIFAR10 0.94 / 0.95 6.3 / 9.6 0.89 / 0.88 6.4 / 7.1
VAE MNIST 0.88 / 0.95 8.5 / 9.3 0.87 / 0.96 2.3 / 3.6
CIFAR10 0.96 / 0.95 8.1 / 8.7 0.95 / 0.93 2.4 / 2.7
VGG16 MNIST 98 / 94 11.9 / 11.8 98 / 97 7.8 / 8.9
CIFAR10 90 / 86 1.9 / 4.0 75 / 89 31.6 / 32.6
GTSRB 85 / 83 4.9 / 6.5 51 / 74 25.8 / 25.8
TIN 55 / 50 2.9 / 3.3 52 / 44 37.8 / 38.6
ResNet18 MNIST 98 / 94 4.5 / 6.7 99 / 98 5.7 / 6.4
CIFAR10 91 / 87 0.8 / 3.0 91 / 91 22.7 / 22.6
GTSRB 93 / 88 3.8 / 3.6 92 / 92 12.5 / 13.6
TIN 56 / 52 1.3 / 1.4 54 / 54 25.3 / 24.8
@@
Figure 10: Degradation of generation performance. Translation images produced with SSIM $\leq$ 0.80 to the regular GAN-generated images. At 0.80 SSIM and below, the generated images have visible issues such as blurriness and background saturation.
V-E3 Fine-tuning with Regularization
Figure 11 contains results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy in percentage for VGG16 and ResNet18 trained on MNIST and CIFAR10, respectively. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, this comes at the cost of accuracy. We observe that a large regularization factor decreases both energy consumption and accuracy. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. As a consequence, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. The results for other models and datasets show the same patterns as Figure 11.
Figure 11: Evaluation of the regularization defense on SkipSponge and Sponge Poisoning for VGG16 and ResNet18 on MNIST and CIFAR10.
V-F User Study
Methodology. We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
Second round. A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and have knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population., and the result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.

### v4

@@ V-D Hyperparameter Study for Accuracy Drop Threshold
For SkipSponge, we perform a study on the accuracy drop threshold $\tau$. In Figures 6 and 12, we show the cumulative energy increase over all the target layers with different accuracy thresholds for VGG16, with $\tau\in\{0\%,1\%,2\%,5\%\}$. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. This is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. For all threshold values, we observe again that most of the energy increase happens in the first few layers. This is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN. We hypothesize that the negligible energy increase in the later layers is partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 7 and 13. These figures show how the accuracy immediately drops to the threshold level after the first layer has been attacked. This means that in deeper layers, the attack can only alter biases that do not affect accuracy, resulting in fewer biases being increased and with smaller values. Thus, deeper layers cause less energy increase. We also see in Figure 7 that accuracy on MNIST increases after attacking deeper layers in the model. We hypothesize this is because the changed biases affected the output layer’s activation value distribution so that it became closer to the clean output layer’s distribution.
Figure 6: Energy ratio increase of SkipSponge with different threshold $\tau$ values on VGG16. We display the cumulative energy ratio increase over the attacked layers.
Figure 7: Accuracy of SkipSponge with different threshold $\tau$ values on VGG16. We display the accuracy of the model after each layer has been attacked.
V-E Hyperparameter Study for Step Size
@@
Figure 8 shows the energy increase for the SkipSponge attack on VGG16 trained on CIFAR10 for different values of the step size $\alpha$. The attacker can get more energy increase by using a smaller step size. However, this comes at the cost of computation time. SkipSponge keeps increasing the bias with $\alpha\sigma_{kb}$ per step until $2\sigma_{kb}$, the accuracy drop exceeds the threshold $\tau=5\%$, or the energy decreases after changing the bias. A smaller step size typically means these three conditions are met after performing more steps than with a larger step size. Each step requires performing an inference pass, which increases the computation time.
@@
Figure 8: Energy ratio increase of SkipSponge with different step size $\alpha$ values on VGG16 trained on CIFAR10. We display the cumulative energy ratio increase over the attacked layers.
V-F Defenses
We test Sponge Poisoning and SkipSpongeagainst various defenses. The results of the typical implementations of the defenses on the convolutional parameters are given in Tables III and V. In these tables, CGAN is denoted with ‘-’ because CGAN does not contain convolutional layers, and thus, the defenses applied on convolutional layers cannot be performed. The left side operand of $\rightarrow$ shows the value before the defense, while the right side operand shows the value after applying the defense. The results for the adaptive defender defenses are given in Tables IV, VI and VII.
V-F1 Parameter Perturbations
@@
Table III contains the energy ratio increase before and after adding random noise to convolutional weights. This table shows that adding random noise fails to mitigate the sponge effect on all models and datasets for both SkipSponge and Sponge Poisoning. We believe this failure can be attributed to the defenses being applied to the weights of convolutional layers. Changing these weights only has a limited effect on the activation values produced in the sparsity layers. Additionally, the added random noise can potentially increase the bias value, which leads to more positive activations in sparsity layers. In contrast, Table IV shows that the adapted defense is more effective at mitigating the sponge effect on all models except the Sponge-Poisoned image classification models. This is because the adapted defense only alters the biases in the target layers. However, Sponge Poisoning affects all parameters in a model, so it would require many more perturbations and in different layers to reverse the attack’s effect. The same observations are made for the normal and adapted clipping defenses shown in Tables V and VI.
@@
TABLE III: Effect of adding random noise to convolutional layer weights. ‘-’ indicates that the value is not applicable.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 5.1 1.5 $\rightarrow$ 1.7
Black hair 5.3 $\rightarrow$ 5.3 1.6 $\rightarrow$ 1.4
CGAN MNIST - -
AE MNIST 13.1 $\rightarrow$ 12.6 4.4 $\rightarrow$ 4.0
CIFAR10 9.6 $\rightarrow$ 9.9 7.1 $\rightarrow$ 6.8
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.5
CIFAR10 8.7 $\rightarrow$ 8.9 2.7 $\rightarrow$ 2.7
VGG16 MNIST 11.8 $\rightarrow$ 11.9 8.9 $\rightarrow$ 8.5
CIFAR10 4.0 $\rightarrow$ 4.1 32.6 $\rightarrow$ 32.4
GTSRB 6.5 $\rightarrow$ 6.7 25.8 $\rightarrow$ 25.8
TIN 3.3 $\rightarrow$ 3.3 38.6 $\rightarrow$ 38.4
ResNet18 MNIST 6.5 $\rightarrow$ 6.7 6.4 $\rightarrow$ 6.0
CIFAR10 3.0 $\rightarrow$ 2.9 22.6 $\rightarrow$ 22.5
GTSRB 3.6 $\rightarrow$ 3.5 13.6 $\rightarrow$ 13.4
TIN 1.4 $\rightarrow$ 1.4 24.8 $\rightarrow$ 25.0
TABLE IV: Effect of adding random noise to target layer biases.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 2.1 1.5 $\rightarrow$ -1.2
Black hair 5.3 $\rightarrow$ 2.4 1.4 $\rightarrow$ - 1.1
CGAN MNIST 4.9 $\rightarrow$ 3.6 0.1 $\rightarrow$ 0.0
AE MNIST 13.1 $\rightarrow$ 12.8 4.4 $\rightarrow$ 3.5
CIFAR10 9.6 $\rightarrow$ 7.3 7.1 $\rightarrow$ 4.7
VAE MNIST 9.3 $\rightarrow$ 9.2 3.6 $\rightarrow$ 3.1
CIFAR10 8.7 $\rightarrow$ 6.0 2.7 $\rightarrow$ 2.2
VGG16 MNIST 11.8 $\rightarrow$ 11.2 8.9 $\rightarrow$ 7.9
CIFAR10 4.0 $\rightarrow$ 0.4 32.6 $\rightarrow$ 32.3
GTSRB 6.5 $\rightarrow$ 3.9 25.8 $\rightarrow$ 25.4
TIN 3.3 $\rightarrow$ 1.8 38.6 $\rightarrow$ 38.5
ResNet18 MNIST 6.7 $\rightarrow$ 4.6 6.4 $\rightarrow$ 5.7
CIFAR10 3.0 $\rightarrow$ 0.8 22.6 $\rightarrow$ 22.6
GTSRB 3.6 $\rightarrow$ 2.0 13.6 $\rightarrow$ 12.9
TIN 1.4 $\rightarrow$ 0.7 24.8 $\rightarrow$ 24.6
@@
TABLE V: Effect of clipping convolutional layer weights. ‘-’ indicates that the value is not applicable.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ 4.6 1.5 $\rightarrow$ 1.5
Black hair 5.3 $\rightarrow$ 5.2 1.4 $\rightarrow$ 1.7
CGAN MNIST - -
AE MNIST 13.1 $\rightarrow$ 12.9 4.4 $\rightarrow$ 3.2
CIFAR10 9.6 $\rightarrow$ 8.0 7.1 $\rightarrow$ 4.6
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 3.9
CIFAR10 8.7 $\rightarrow$ 5.6 2.7 $\rightarrow$ 2.3
VGG16 MNIST 11.8 $\rightarrow$ 12.9 8.9 $\rightarrow$ 10.4
CIFAR10 4.0 $\rightarrow$ 4.2 32.6 $\rightarrow$ 32.9
GTSRB 6.5 $\rightarrow$ 6.2 25.8 $\rightarrow$ 26.0
TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.7
ResNet18 MNIST 6.7 $\rightarrow$ 6.8 6.4 $\rightarrow$ 6.5
CIFAR10 3.0 $\rightarrow$ 3.1 22.6 $\rightarrow$ 22.6
GTSRB 3.6 $\rightarrow$ 3.7 13.6 $\rightarrow$ 14.1
TIN 1.4 $\rightarrow$ 1.5 24.8 $\rightarrow$ 25.0
TABLE VI: Effect of clipping target layer biases.
Model Dataset SkipSponge SP
StarGAN Age 4.8 $\rightarrow$ -2.2 1.5 $\rightarrow$ -1.2
Black hair 5.3 $\rightarrow$ -1.9 1.4 $\rightarrow$ -0.9
CGAN MNIST 4.9 $\rightarrow$ 3.1 0.1 $\rightarrow$ -0.1
AE MNIST 13.1 $\rightarrow$ 12.5 4.4 $\rightarrow$ 2.9
CIFAR10 9.6 $\rightarrow$ 8.7 7.1 $\rightarrow$ 6.5
VAE MNIST 9.3 $\rightarrow$ 9.4 3.6 $\rightarrow$ 2.6
CIFAR10 8.7 $\rightarrow$ 7.5 2.7 $\rightarrow$ 1.5
VGG16 MNIST 11.8 $\rightarrow$ 10.7 8.9 $\rightarrow$ 0.1
CIFAR10 4.0 $\rightarrow$ 1.5 32.6 $\rightarrow$ 32.3
GTSRB 6.5 $\rightarrow$ 6.1 25.8 $\rightarrow$ 25.9
TIN 3.3 $\rightarrow$ 3.2 38.6 $\rightarrow$ 38.5
ResNet18 MNIST 6.7 $\rightarrow$ 6.1 6.4 $\rightarrow$ 4.1
CIFAR10 3.0 $\rightarrow$ - 0.1 22.6 $\rightarrow$ 22.7
GTSRB 3.6 $\rightarrow$ - 1.3 13.6 $\rightarrow$ 13.0
TIN 1.4 $\rightarrow$ 1.0 24.8 $\rightarrow$ 24.9
V-F2 Fine-pruning
@@
Table VII contains the energy ratio increase and the accuracy or SSIM after applying the adapted fine-pruning defense. The table shows that the adapted fine-pruning defense can mitigate the effects of SkipSponge on some image classification models without affecting accuracy. Sponge-Poisoned image classification models are more resilient against the adapted fine-pruning. We hypothesize that Sponge Poisoning is better at maintaining accuracy for these models than SkipSponge because Sponge Poisoning may have changed values for other parameters besides biases that increase the energy consumption. Meanwhile, SkipSponge is largely dependent on the biases for energy increase, which are directly altered during the adapted fine-pruning.
@@
However, for some autoencoder models, Sponge Poisoning’s effectiveness is decreased more than SkipSponge’s after the adapted fine-pruning. Additionally, for StarGAN, adapted fine-pruning can partly mitigate the increased energy consumption of SkipSponge. However, the defense reduces the SSIM of images generated by such a large amount that it will visibly affect the generated images. This can be seen in Figure 9. The figure contains images generated by SkipSponge StarGAN for 0.8 SSIM. Sponge-poisoned StarGAN shows similar results at 0.8 SSIM. The images show how an SSIM at and below 0.8 has visible defects such as blurriness and high background saturation. Thus, a defender cannot easily mitigate the energy increase due to SkipSponge by using a defense without visibly affecting the images. Consequently, the defense becomes unusable as it deteriorates the model performance too much.
@@
TABLE VII: Effect of fine-pruning on target layer biases. The Acc. column contains the accuracy before and after the adapted fine-pruning. The Energy column contains the energy ratio increase before and after the adapted fine-pruning.
Model Dataset SkipSponge SP
@@
StarGAN Age 95 $\rightarrow$ 84 4.8 $\rightarrow$ 1.3 84 $\rightarrow$ 83 1.5 $\rightarrow$ 1.0
Black hair 95 $\rightarrow$ 80 5.3 $\rightarrow$ 1.2 84 $\rightarrow$ 81 1.4 $\rightarrow$ 0.9
CGAN MNIST 95 $\rightarrow$ 89 4.9 $\rightarrow$ 3.6 49 $\rightarrow$ 51 0.1 $\rightarrow$ 0.1
AE MNIST 95 $\rightarrow$ 96 13.1 $\rightarrow$ 10.2 93 $\rightarrow$ 94 4.4 $\rightarrow$ 1.8
CIFAR10 95 $\rightarrow$ 94 9.6 $\rightarrow$ 6.3 88 $\rightarrow$ 89 7.1 $\rightarrow$ 6.4
VAE MNIST 95 $\rightarrow$ 88 9.3 $\rightarrow$ 8.5 96 $\rightarrow$ 87 3.6 $\rightarrow$ 2.3
CIFAR10 95 $\rightarrow$ 96 8.7 $\rightarrow$ 8.1 93 $\rightarrow$ 95 2.7 $\rightarrow$ 2.4
VGG16 MNIST 94 $\rightarrow$ 98 11.8 $\rightarrow$ 11.9 97 $\rightarrow$ 98 8.9 $\rightarrow$ 7.8
CIFAR10 86 $\rightarrow$ 90 4.0 $\rightarrow$ 1.9 89 $\rightarrow$ 75 32.6 $\rightarrow$ 31.6
GTSRB 83 $\rightarrow$ 85 6.5 $\rightarrow$ 4.9 74 $\rightarrow$ 51 25.8 $\rightarrow$ 25.8
TIN 55 $\rightarrow$ 55 3.3 $\rightarrow$ 2.9 44 $\rightarrow$ 52 38.6 $\rightarrow$ 37.8
ResNet18 MNIST 94 $\rightarrow$ 98 6.7 $\rightarrow$ 4.5 98 $\rightarrow$ 99 6.4 $\rightarrow$ 5.7
CIFAR10 87 $\rightarrow$ 91 3.0 $\rightarrow$ 0.8 91 $\rightarrow$ 91 22.6 $\rightarrow$ 22.7
GTSRB 88 $\rightarrow$ 93 3.6 $\rightarrow$ 3.8 92 $\rightarrow$ 92 13.6 $\rightarrow$ 12.5
TIN 52 $\rightarrow$ 56 1.4 $\rightarrow$ 1.3 54 $\rightarrow$ 54 24.8 $\rightarrow$ 25.3
@@
Figure 9: Images generated with an SSIM $\leq$ 0.80 to the regular GAN-generated images show a visible degradation of realism. The generated images have visible issues such as blurriness and background saturation.
V-F3 Fine-tuning with Regularization
Figure 10 contains the results of the fine-tuning with regularization defense experiments. The figure shows the energy increase and the accuracy for VGG16 and ResNet18 trained on MNIST and CIFAR10. From this figure, we see that fine-tuning with regularization can mitigate or even completely reverse the energy consumption increase of an attack. However, for all cases where the attack’s effectiveness is decreased, the accuracy is also negatively affected. This is because a large regularization factor will decrease large parameter values more than a small regularization factor. Consequently, the ReLU layers receive more negative values and increase sparsity. Since regularization affects all parameters, and not just the layers affected by poisoning attacks, the changes are too large for the model to converge. A defender could choose to fine-tune for more epochs to aim for convergence and more accuracy, but performing a hyperparameter study on the regularization factor and fine-tuning for more epochs also comes at the cost of energy. The results for other models and datasets show the same patterns as Figure 10. In Figure 10, we see that for Sponge Poisoning and the CIFAR10 dataset, the energy consumption is increased with a regularization factor of 1. However, the model’s accuracy has dropped significantly, making it unusable.
Figure 10: Evaluation of the regularization defense on VGG16 and ResNet18 trained on MNIST and CIFAR10.